\begin{thebibliography}{10}

\bibitem{key-1}
M.t A{\c{c}}ikg{\"o}z, D.~Erdal, and S.~Araci, \emph{A new approach to
  {$q$}-{B}ernoulli numbers and {$q$}-{B}ernoulli polynomials related to
  {$q$}-{B}ernstein polynomials}, Adv. Difference Equ. (2010), Art. ID 951764,
  9. \MR{2764538 (2012g:11038)}

\bibitem{key-2}
S.~Araci, M.~Acikgoz, and H.~Jolany, \emph{On the families of {$q$}-{E}uler
  polynomials and their applications}, J. Egyptian Math. Soc. \textbf{23}
  (2015), no.~1, 1--5. \MR{3317290}

\bibitem{key-4}
A.~Bayad and S.~Gaboury, \emph{Generalized {D}irichlet {$L$}-function of
  arbitrary order with applications}, Adv. Stud. Contemp. Math. (Kyungshang)
  \textbf{23} (2013), no.~4, 607--619. \MR{3115219}

\bibitem{key-3}
A.~Bayad and Y.~Hamahata, \emph{Multiple polylogarithms and
  multi-poly-{B}ernoulli polynomials}, Funct. Approx. Comment. Math.
  \textbf{46} (2012), no.~part 1, 45--61. \MR{2951728}

\bibitem{key-7}
L.~Carlitz, \emph{A degenerate {S}taudt-{C}lausen theorem}, Arch. Math. (Basel)
  \textbf{7} (1956), 28--33. \MR{0074436 (17,586a)}

\bibitem{key-6}
\bysame, \emph{Degenerate {S}tirling, {B}ernoulli and {E}ulerian numbers},
  Utilitas Math. \textbf{15} (1979), 51--88. \MR{531621 (80i:05014)}

\bibitem{key-8}
R.~Dere and Y.~Simsek, \emph{Applications of umbral algebra to some special
  polynomials}, Adv. Stud. Contemp. Math. (Kyungshang) \textbf{22} (2012),
  no.~3, 433--438. \MR{2976601}

\bibitem{key-9}
D.~Ding and J.~Yang, \emph{Some identities related to the {A}postol-{E}uler and
  {A}postol-{B}ernoulli polynomials}, Adv. Stud. Contemp. Math. (Kyungshang)
  \textbf{20} (2010), no.~1, 7--21. \MR{2597988 (2011k:05030)}

\bibitem{key-5}
S.~Gaboury, R.~Tremblay, and B.-J. Fug{\`e}re, \emph{Some explicit formulas for
  certain new classes of {B}ernoulli, {E}uler and {G}enocchi polynomials},
  Proc. Jangjeon Math. Soc. \textbf{17} (2014), no.~1, 115--123. \MR{3184467}

\bibitem{key-10}
H.~Jolany and H.~Mohebbi, \emph{Some results on {G}eneralized multi
  poly-{B}ernoulli and {E}uler polynomials}, Math. Combin. Book Ser. \textbf{2}
  (2011), 117--129.

\bibitem{key-15}
D.~S. Kim and T.~Kim, \emph{Hermite and poly-{B}ernoulli mixed-type
  polynomials}, Adv. Difference Equ. (2013), 2013:343, 12. \MR{3213908}

\bibitem{key-14}
\bysame, \emph{Higher-order {F}robenius-{E}uler and poly-{B}ernoulli mixed-type
  polynomials}, Adv. Difference Equ. (2013), 2013:251. \MR{3108262}

\bibitem{key-16}
\bysame, \emph{A note on poly-{B}ernoulli and higher-order poly-{B}ernoulli
  polynomials}, Russ. J. Math. Phys. \textbf{22} (2015), no.~1, 26--33.
  \MR{3318873}

\bibitem{key-13}
\bysame, \emph{Some identities of degenerate {E}uler polynomials arising from
  {$p$}-adic fermionic integrals on {$\Bbb{Z}_p$}}, Integral Transforms Spec.
  Funct. \textbf{26} (2015), no.~4, 295--302. \MR{3304355}

\bibitem{key-12}
D.~S. Kim, T.~Kim, D.~V. Dolgy, and T.~Komatsu, \emph{Barnes-type degenerate
  {B}ernoulli polynomials}, Adv. Stud. Contemp. Math. \textbf{25} (2015),
  no.~1, 121--146.

\bibitem{key-17}
D.~S. Kim, T.~Kim, and S.~H. Lee, \emph{A note on poly-bernoulli polynomials
  arising umbral calculus}, Adv. Studies Theor. Phys. \textbf{7} (2013),
  no.~15, 731--744.

\bibitem{key-11}
D.~S. Kim, T.~Kim, T.~Mansour, and D.~V. Dolgy, \emph{On {P}oly-{B}ernoulli
  polynomials of the second kind with umbral calculus viewpoint}, Adv.
  Difference Equ. (2015), no.~1, 2015:27. \MR{3304954}

\bibitem{key-21}
T.~Kim, \emph{Some identities on the {$q$}-{E}uler polynomials of higher order
  and {$q$}-{S}tirling numbers by the fermionic {$p$}-adic integral on {$\Bbb
  Z_p$}}, Russ. J. Math. Phys. \textbf{16} (2009), no.~4, 484--491. \MR{2587805
  (2011e:33045)}

\bibitem{key-19}
\bysame, \emph{Barnes' type multiple degenerate {B}ernoulli and {E}uler
  polynomials}, Applied Math. Comput. \textbf{258} (2015), 556--564.

\bibitem{key-18}
T.~Kim, Y.~S. Jang, and J.~J. Seo, \emph{poly-{B}ernoulli polynomials and their
  applications}, Int. Journal of Math. Analysis \textbf{8} (2014), no.~30,
  1495--1503.

\bibitem{key-20}
T.~Kim, H.~K. Kwon, S.~H. Lee, and J.~J. Seo, \emph{A note on poly-{B}ernoulli
  numbers and polynomials of the second kind}, Adv. Difference Equ. (2014),
  2014:219.

\bibitem{key-22}
Q.-M. Luo, B.-N. Guo, and F.~Qi, \emph{On evaluation of {R}iemann zeta function
  {$\zeta(s)$}}, Adv. Stud. Contemp. Math. (Kyungshang) \textbf{7} (2003),
  no.~2, 135--144. \MR{2005861 (2004h:11069)}

\bibitem{key-23}
Q.-M. Luo and F.~Qi, \emph{Relationships between generalized {B}ernoulli
  numbers and polynomials and generalized {E}uler numbers and polynomials},
  Adv. Stud. Contemp. Math. (Kyungshang) \textbf{7} (2003), no.~1, 11--18.
  \MR{1981601}

\bibitem{key-24}
J.-W. Park and S.-H. Rim, \emph{On the modified {$q$}-{B}ernoulli polynomials
  with weight}, Proc. Jangjeon Math. Soc. \textbf{17} (2014), no.~2, 231--236.
  \MR{3234991}

\bibitem{key-25}
E.~{\c{S}}en, \emph{Theorems on {A}postol-{E}uler polynomials of higher order
  arising from {E}uler basis}, Adv. Stud. Contemp. Math. (Kyungshang)
  \textbf{23} (2013), no.~2, 337--345. \MR{3088764}

\bibitem{key-26}
H.~M. Srivastava, T.~Kim, and Y.~Simsek, \emph{{$q$}-{B}ernoulli numbers and
  polynomials associated with multiple {$q$}-zeta functions and basic
  {$L$}-series}, Russ. J. Math. Phys. \textbf{12} (2005), no.~2, 241--268.
  \MR{2199005 (2006k:11034)}

\bibitem{key-27}
Z.~Zhang and J.~Yang, \emph{On sums of products of the degenerate {B}ernoulli
  numbers}, Integral Transforms Spec. Funct. \textbf{20} (2009), no.~9-10,
  751--755.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amsthm}
\usepackage{amstext}
\usepackage{amssymb}
\usepackage{hyperref}

\makeatletter
\numberwithin{equation}{section}
\numberwithin{figure}{section}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
  \theoremstyle{remark}
  \newtheorem*{rem*}{Remark}

\allowdisplaybreaks
\usepackage[sort&compress,square,numbers]{natbib}

\newcommand\relphantom[1]{\mathrel{\phantom{#1}}} 
\global\long\def\Li{\mathrm{Li}}

\makeatother

\begin{document}

\title{A note on degenerate poly-Bernoulli numbers and polynomials}

\author{Dae San Kim}
\address{Department of Mathematics, Sogang University, Seoul 121-742, Republic
of Korea}
\email{dskim@sogang.ac.kr}

\author{Taekyun Kim}
\address{Department of Mathematics, Kwangwoon University, Seoul 139-701, Republic
of Korea}
\email{tkkim@kw.ac.kr}

\keywords{degenerate poly-Bernoulli polynomial, degenerate Bernoulli polynomial, Stirling number of the second kind }
\subjclass[2010]{11B68, 11B73, 11B83}

\begin{abstract}
In this paper, we consider the degenerate poly-Bernoulli polynomials
and present new and explicit formulas for computing them in terms of the degenerate Bernoulli polynomials and Stirling numbers of the second kind.
\end{abstract}

\maketitle

\section{Introduction}

For $\lambda\in\mathbb{C}$, L. Carlitz considered the degenerate
Bernoulli polynomials given by the generating function 
\begin{equation}
\frac{t}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}=\sum_{n=0}^{\infty}\beta_{n}\left(x\mid\lambda\right)\frac{t^{n}}{n!},\quad\left(\text{see \cite{key-6,key-12,key-19}}\right).\label{eq:1}
\end{equation}

When $x=0$, $\beta_{n}\left(\lambda\right)=\beta_{n}\left(0\mid\lambda\right)$
are called the degenerate Bernoulli numbers. 

Thus, by (\ref{eq:1}), we get 
\begin{equation}
\beta_{n}\left(x\mid\lambda\right)=\sum_{l=0}^{n}\binom{n}{l}\beta_{l}\left(\lambda\right)\left(x\mid\lambda\right)_{n-l},\label{eq:2}
\end{equation}
where $\left(x\mid\lambda\right)_{n}=x\left(x-\lambda\right)\left(x-2\lambda\right)\cdots\left(x-\lambda\left(n-1\right)\right)$. 

The classical polylograithm function $\Li_{k}$ is 
\begin{equation}
\Li_{k}\left(x\right)=\sum_{n=1}^{\infty}\frac{x^{n}}{n^{k}},\quad\left(k\in\mathbb{Z}\right),\quad\left(\text{see \cite{key-10,key-11,key-12,key-13,key-14,key-15,key-16,key-17,key-18}}\right).\label{eq:3}
\end{equation}

From (\ref{eq:1}), we note that 
\begin{align}
 & \relphantom{=}{}\sum_{n=0}^{\infty}\lim_{\lambda\rightarrow0}\beta_{n}\left(x\mid\lambda\right)\frac{t^{n}}{n!}\label{eq:4}\\
 & =\lim_{\lambda\rightarrow0}\frac{t}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\nonumber \\
 & =\frac{t}{e^{t}-1}e^{xt}\nonumber \\
 & =\sum_{n=0}^{\infty}B_{n}\left(x\right)\frac{t^{n}}{n!},\nonumber 
\end{align}
where $B_{n}\left(x\right)$ are called the Bernoulli polynomials
(see \cite{key-1,key-2,key-3,key-4,key-5,key-6,key-7,key-8,key-9,key-10,key-11,key-12,key-13,key-14,key-15,key-16,key-17,key-18,key-19,key-20,key-21,key-22,key-23,key-24,key-25,key-26,key-27}). 

Thus, by (\ref{eq:4}), we get 
\begin{equation}
\lim_{\lambda\rightarrow0}\beta_{n}\left(x\mid\lambda\right)=B_{n}\left(x\right),\quad\left(n\ge0\right).\label{eq:5}
\end{equation}

In \cite{key-3,key-10}, the poly-Bernoulli polynomials are given
by 
\begin{equation}
\frac{\Li_{k}\left(1-e^{-t}\right)}{e^{t}-1}e^{xt}=\sum_{n=0}^{\infty}B_{n}^{\left(k\right)}\left(x\right)\frac{t^{n}}{n!}.\label{eq:6}
\end{equation}

For $k=1$, we have 
\begin{equation}
\frac{\Li_{1}\left(1-e^{-t}\right)}{e^{t}-1}e^{xt}=\frac{t}{e^{t}-1}e^{xt}=\sum_{n=0}^{\infty}B_{n}\left(x\right)\frac{t^{n}}{n!}.\label{eq:7}
\end{equation}

By (\ref{eq:4}) and (\ref{eq:7}), we get $B_{n}^{\left(1\right)}\left(x\right)=B_{n}\left(x\right)$. 

The Stirling numbers of the second kind are given by 
\begin{equation}
x^{n}=\sum_{l=0}^{n}S_{2}\left(n,l\right)\left(x\right)_{l},\quad\left(\text{see \cite{key-1,key-2,key-3,key-4,key-5,key-6,key-7,key-8,key-9,key-10,key-11,key-12,key-13,key-14,key-15,key-16,key-17,key-18,key-19,key-20,key-21,key-22,key-23,key-24,key-25,key-26,key-27}}\right).\label{eq:8}
\end{equation}
and the Stirling numbers of the first kind are defined by 
\begin{equation}
\left(x\right)_{n}=x\left(x-1\right)\cdots\left(x-n+1\right)=\sum_{l=0}^{n}S_{1}\left(n,l\right)x^{l},\quad\left(n\ge0\right).\label{eq:9}
\end{equation}

The purpose of this paper is to construct the degenerate poly-Bernoulli
polynomials and present new and explicit formulas for computing them
in terms of the degenerate Bernoulli polynomials and Stirling
numbers of the second kind.

\section{Degenerate poly-Bernoulli numbers and polynomials}

For $\lambda\in\mathbb{C}$, $k\in\mathbb{Z}$, we consider the degenerate
poly-Bernoulli polynomials given by the generating function

\begin{equation}
\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}=\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}.\label{eq:10}
\end{equation}

When $x=0$, $\beta_{n}^{\left(k\right)}\left(\lambda\right)=\beta_{n}^{\left(k\right)}\left(0\mid\lambda\right)$
are called the degenerate poly-Bernoulli numbers. Note that $\beta_{n}^{\left(1\right)}\left(x\mid\lambda\right)=\beta_{n}\left(x\mid\lambda\right)$
and $\lim_{\lambda\rightarrow0}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)=B_{n}^{\left(k\right)}\left(x\right)$. 

From (\ref{eq:10}), we can derive the following equation: 
\begin{align}
\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!} & =\left(\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\right)\left(1+\lambda t\right)^{\frac{x}{\lambda}}\label{eq:11}\\
 & =\left(\sum_{l=0}^{\infty}\beta_{l}^{\left(k\right)}\left(\lambda\right)\frac{t^{l}}{l!}\right)\left(\sum_{m=0}^{\infty}\left(x\mid\lambda\right)_{m}\frac{t^{m}}{m!}\right)\nonumber \\
 & =\sum_{n=0}^{\infty}\left(\sum_{l=0}^{n}\binom{n}{l}\beta_{l}^{\left(k\right)}\left(\lambda\right)\left(x\mid\lambda\right)_{n-l}\right)\frac{t^{n}}{n!}.\nonumber 
\end{align}

Thus, by (\ref{eq:11}), we get 
\begin{equation}
\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)=\sum_{l=0}^{n}\binom{n}{l}\beta_{l}^{\left(k\right)}\left(\lambda\right)\left(x\mid\lambda\right)_{n-l}.\label{eq:12}
\end{equation}

Now, we observe that 
\begin{align}
 & \relphantom{=}{}\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+t\right)^{\frac{x}{\lambda}}\label{eq:13}\\
 & =\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}\nonumber \\
 & =\frac{\left(1+t\right)^{\frac{x}{\lambda}}}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\underset{\left(k-2\right)\text{times}}{\int_{0}^{t}\underbrace{\frac{1}{e^{y}-1}\int_{0}^{y}\frac{1}{e^{y}-1}\int_{0}^{y}\cdots\frac{1}{e^{y}-1}\int_{0}^{y}}}\frac{y}{e^{y}-1}dy\cdots dy.\nonumber 
\end{align}

From (\ref{eq:13}), we have 
\begin{align}
 & \relphantom{=}{}\sum_{n=0}^{\infty}\beta_{n}^{\left(2\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}\label{eq:14}\\
 & =\frac{\left(1+t\right)^{\frac{x}{\lambda}}}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\int_{0}^{t}\frac{y}{e^{y}-1}dy\nonumber \\
 & =\frac{\left(1+t\right)^{\frac{x}{\lambda}}}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\sum_{l=0}^{\infty}\frac{B_{l}}{l!}\int_{0}^{t}y^{l}dy\nonumber \\
 & =\left(\frac{t}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\right)\left(\sum_{l=0}^{\infty}\frac{B_{l}}{l+1}\frac{t^{l}}{l!}\right)\nonumber \\
 & =\sum_{n=0}^{\infty}\left\{ \sum_{l=0}^{n}\binom{n}{l}\frac{B_{l}}{l+1}\beta_{n-l}\left(x\mid\lambda\right)\right\} \frac{t^{n}}{n!},\nonumber 
\end{align}
where $B_{n}=B_{n}\left(0\right)$ are Bernoulli numbers.

By comparing the coefficients on both sides of (\ref{eq:14}),
we obtain the following theorem.
\begin{thm}
\label{thm:1} For $n\ge0$, we have 
\begin{align*}
\beta_{n}^{\left(2\right)}\left(x\mid\lambda\right) & =\sum_{l=0}^{n}\binom{n}{l}\frac{B_{l}}{l+1}\beta_{n-l}\left(x\mid\lambda\right)\\
 & =\beta_{n}\left(x\mid\lambda\right)-\frac{n}{4}\beta_{n-1}\left(x\mid\lambda\right)+\sum_{l=2}^{n}\binom{n}{l}\frac{B_{l}}{l+1}\beta_{n-l}\left(x\mid\lambda\right).
\end{align*}

Moreover, 
\[
\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)=\sum_{l=0}^{n}\binom{n}{l}\beta_{l}^{\left(k\right)}\left(\lambda\right)\left(x\mid\lambda\right)_{n-l}.
\]

\end{thm}
By (\ref{eq:13}), we easily get
\begin{align}
 & \relphantom{=}{}\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}\label{eq:15}\\
 & =\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+t\right)^{\frac{x}{\lambda}}\nonumber \\
 & =\frac{t}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+t\right)^{\frac{x}{\lambda}}\frac{\Li_{k}\left(1-e^{-t}\right)}{t}.\nonumber 
\end{align}

We observe that 

\begin{align}
\frac{1}{t}\Li_{k}\left(1-e^{-t}\right) & =\frac{1}{t}\sum_{n=1}^{\infty}\frac{1}{n^{k}}\left(1-e^{-t}\right)^{n}\label{eq:16}\\
 & =\frac{1}{t}\sum_{n=1}^{\infty}\frac{\left(-1\right)^{n}}{n^{k}}n!\sum_{l=n}^{\infty}S_{2}\left(l,n\right)\frac{\left(-t\right)^{l}}{l!}\nonumber \\
 & =\frac{1}{t}\sum_{l=1}^{\infty}\sum_{n=1}^{l}\frac{\left(-1\right)^{n+l}}{n^{k}}n!S_{2}\left(l,n\right)\frac{t^{l}}{l!}\nonumber \\
 & =\sum_{l=0}^{\infty}\sum_{n=1}^{l+1}\frac{\left(-1\right)^{n+l+1}}{n^{k}}n!\frac{S_{2}\left(l+1,n\right)}{l+1}\frac{t^{l}}{l!}.\nonumber 
\end{align}

From (\ref{eq:15}) and (\ref{eq:16}), we have 
\begin{align}
 & \relphantom{=}{}\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}\label{eq:17}\\
 & =\left(\sum_{m=0}^{\infty}\beta_{m}\left(x\mid\lambda\right)\frac{t^{m}}{m!}\right)\left(\sum_{l=0}^{\infty}\left(\sum_{p=1}^{l+1}\frac{\left(-1\right)^{p+l+1}}{p^{k}}p!\frac{S_{2}\left(l+1,p\right)}{l+1}\right)\frac{t^{l}}{l!}\right)\nonumber \\
 & =\sum_{n=0}^{\infty}\left\{ \sum_{l=0}^{n}\binom{n}{l}\left(\sum_{p=1}^{l+1}\frac{\left(-1\right)^{p+l+1}p!}{p^{k}}\frac{S_{2}\left(l+1,p\right)}{l+1}\right)\beta_{n-l}\left(x\mid\lambda\right)\right\} \frac{t^{n}}{n!}.\nonumber 
\end{align}

By comparing the coefficients on both sides of (\ref{eq:17}),
we obtain the following theorem.
\begin{thm}
\label{thm:2} For $n\ge0$, we have 
\[
\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)=\sum_{l=0}^{n}\binom{n}{l}\left(\sum_{p=1}^{l+1}\frac{\left(-1\right)^{p+l+1}p!}{p^{k}}\frac{S_{2}\left(l+1,p\right)}{l+1}\right)\beta_{n-l}\left(x\mid\lambda\right).
\]

\end{thm}
It is easy to show that 
\begin{align}
 & \relphantom{=}\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x+1}{\lambda}}-\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\label{eq:18}\\
 & =\left(1+\lambda t\right)^{\frac{x}{\lambda}}\Li_{k}\left(1-e^{-t}\right)\nonumber \\
 & =\left(\sum_{l=0}^{\infty}\left(x\mid\lambda\right)_{l}\frac{t^{l}}{l!}\right)\left(\sum_{m=1}^{\infty}\frac{\left(1-e^{-t}\right)^{m}}{m^{k}}\right)\nonumber \\
 & =\left(\sum_{l=0}^{\infty}\left(x\mid\lambda\right)_{l}\frac{t^{l}}{l!}\right)\left(\sum_{m=0}^{\infty}\frac{\left(1-e^{-t}\right)^{m+1}}{\left(m+1\right)^{k}}\right)\nonumber \\
 & =\left(\sum_{l=0}^{\infty}\left(x\mid\lambda\right)_{l}\frac{t^{l}}{l!}\right)\left(\sum_{p=1}^{\infty}\left(\sum_{m=0}^{p-1}\frac{\left(-1\right)^{m+p+1}}{\left(m+1\right)^{k}}\left(m+1\right)!S_{2}\left(p,m+1\right)\right)\frac{t^{p}}{p!}\right)\nonumber \\
 & =\sum_{n=1}^{\infty}\left\{ \sum_{p=1}^{n}\sum_{m=0}^{p-1}\frac{\left(-1\right)^{m+p+1}}{\left(m+1\right)^{k}}\left(m+1\right)!S_{2}\left(p,m+1\right)\binom{n}{p}\left(x\mid\lambda\right)_{n-p}\right\} \frac{t^{n}}{n!}.\nonumber 
\end{align}

On the other hand, 
\begin{align}
 & \relphantom{=}{}\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x+1}{\lambda}}-\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\label{eq:19}\\
 & =\sum_{n=0}^{\infty}\left\{ \beta_{n}^{\left(k\right)}\left(x+1\mid\lambda\right)-\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\right\} \frac{t^{n}}{n!}.\nonumber 
\end{align}

Therefore, by (\ref{eq:18}) and (\ref{eq:19}), we obtain the following
theorem.
\begin{thm}
\label{thm:3} For $n\ge1$, we have 
\begin{align*}
 & \relphantom{=}{}\beta_{n}^{\left(k\right)}\left(x+1\mid\lambda\right)-\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\\
 & =\sum_{p=1}^{n}\left(\sum_{m=0}^{p-1}\frac{\left(-1\right)^{m+k+1}}{\left(m+1\right)^{k}}\left(m+1\right)!S_{2}\left(k+m+1\right)\right)\binom{n}{p}\left(x\mid\lambda\right)_{n-p}.
\end{align*}

\end{thm}
Now, we note that 
\begin{align}
 & \relphantom{=}{}\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\label{eq:20}\\
 & =\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{d}{\lambda}}-1}\sum_{a=0}^{d-1}\left(1+\lambda t\right)^{\frac{l+x}{\lambda}}\nonumber \\
 & =\left(\frac{\Li_{k}\left(1-e^{-t}\right)}{t}\right)\frac{1}{d}\sum_{a=0}^{d-1}\frac{dt}{\left(1+\lambda t\right)^{\frac{d}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{l+x}{\lambda}}\nonumber\\
 & =\sum_{l=0}^{\infty}\left(\sum_{p=1}^{l+1}\frac{\left(-1\right)^{p+l+1}}{p^{k}}p!\frac{S_{2}\left(l+1,p\right)}{l+1}\right)\frac{t^{l}}{l!}\nonumber \\
 & \relphantom{=}{}\times\sum_{a=0}^{d-1}\sum_{m=0}^{\infty}\beta_{m}\left(\left.\frac{l+x}{d}\right|\frac{\lambda}{d}\right)d^{m-1}\frac{t^{m}}{m!}\nonumber \\
 & =\sum_{a=0}^{d-1}\left(\sum_{n=0}^{\infty}\left(\sum_{l=0}^{n}\sum_{p=1}^{l+1}\binom{n}{l}\frac{\left(-1\right)^{p+l+1}}{p^{k}}p!\frac{S_{2}\left(l+1,p\right)}{l+1}\beta_{n-l}\left(\left.\frac{l+x}{d}\right|\frac{\lambda}{d}\right)d^{n-l-1}\right)\frac{t^{n}}{n!}\right)\nonumber \\
 & =\sum_{n=0}^{\infty}\left\{ \sum_{a=0}^{d-1}\sum_{l=0}^{n}\sum_{p=1}^{l+1}\binom{n}{l}\frac{\left(-1\right)^{p+l+1}}{p^{k}}p!\frac{S_{2}\left(l+1,p\right)}{l+1}\beta_{n-l}\left(\left.\frac{l+x}{d}\right|\frac{\lambda}{d}\right)d^{n-l-1}\right\} \frac{t^{n}}{n!},\nonumber 
\end{align}
where $d$ is a fixed positive integer. 

On the other hand, 
\begin{align}
 & \relphantom{=}{}\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x}{\lambda}}\label{eq:21}\\
 & =\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{n}}{n!}.\nonumber 
\end{align}

Therefore, by (\ref{eq:20}) and (\ref{eq:21}), we obtain the following
theorem.
\begin{thm}
\label{thm:4} For $n\ge0$, $d\in\mathbb{N}$ and $k\in\mathbb{Z}$,
we have 
\begin{align*}
 & \relphantom{=}{}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\\
 & =\sum_{a=0}^{d-1}\sum_{l=0}^{n}\sum_{p=1}^{l+1}\binom{n}{l}\frac{\left(-1\right)^{p+l+1}}{p^{k}}p!\frac{S_{2}\left(l+1,p\right)}{l+1}\beta_{n-l}\left(\left.\frac{l+x}{d}\right|\frac{\lambda}{d}\right)d^{n-l-1}.
\end{align*}

\end{thm}
From (\ref{eq:13}), we can derive the following equation: 
\begin{align}
 & \relphantom{=}{}\sum_{n=0}^{\infty}\beta_{n}^{\left(k\right)}\left(x+y\mid\lambda\right)\frac{t^{n}}{n!}\label{eq:22}\\
 & =\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+\lambda t\right)^{\frac{x+y}{\lambda}}\nonumber \\
 & =\left(\frac{\Li_{k}\left(1-e^{-t}\right)}{\left(1+\lambda t\right)^{\frac{1}{\lambda}}-1}\left(1+t\lambda\right)^{\frac{x}{\lambda}}\right)\left(1+\lambda t\right)^{\frac{y}{\lambda}}\nonumber \\
 & =\left(\sum_{l=0}^{\infty}\beta_{l}^{\left(k\right)}\left(x\mid\lambda\right)\frac{t^{l}}{l!}\right)\left(\sum_{m=0}^{\infty}\left(y\mid\lambda\right)_{m}\frac{t^{m}}{m!}\right)\nonumber \\
 & =\sum_{n=0}^{\infty}\left(\sum_{l=0}^{n}\binom{n}{l}\beta_{l}^{\left(k\right)}\left(x\mid\lambda\right)\left(y\mid\lambda\right)_{n-l}\right)\frac{t^{n}}{n!}.\nonumber 
\end{align}

Therefore, by (\ref{eq:22}), we obtain the following theorem.
\begin{thm}
\label{thm:5} For $n\ge0$, we have 
\[
\beta_{n}^{\left(k\right)}\left(x+y\mid\lambda\right)=\sum_{l=0}^{n}\binom{n}{l}\beta_{l}^{\left(k\right)}\left(x\mid\lambda\right)\left(y\mid\lambda\right)_{n-l}.
\]
\end{thm}
\begin{rem*}
$\,$
\end{rem*}
\begin{align*}
 & \relphantom{=}\frac{d}{dx}\beta_{n}^{\left(k\right)}\left(x\mid\lambda\right)\\
 & =\frac{d}{dx}\sum_{l=0}^{n}\binom{n}{l}\beta_{n-l}^{\left(k\right)}\left(\lambda\right)\left(x\mid\lambda\right)_{l}\\
 & =\sum_{l=0}^{n}\binom{n}{l}\beta_{n-l}^{\left(k\right)}\left(\lambda\right)\sum_{j=0}^{l-1}\frac{1}{x-\lambda j}\prod_{i=0}^{l-1}\left(x-\lambda i\right)\\
 & =\sum_{l=0}^{n}\binom{n}{l}\beta_{n-l}^{\left(k\right)}\left(\lambda\right)\sum_{j=0}^{l-1}\prod_{\substack{i=0\\
i\neq j
}
}^{l-1}\left(x-\lambda i\right).
\end{align*}

{\bf Competing interests}\\
The authors declare that they have no competing interests.\\\\

{\bf Authors' contributions}\\
All authors contributed equally to this work. All authors read and approved the final manuscript.\\\\

\bibliographystyle{amsplain}

\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}